Write a system of two equations in two unknowns for each problem. Solve each system by substitution.
Investing her bonus. Donna invested her $33000 bonus and received a total of $970 in interest after on year. If part of the money returned 4% and the remainder 2.25% the how much did she invest at each rate?

Answer by checkley75(3666)

.04*X+.0225(33000-X)=970
.04X+742.5-.0225X=970
.015X=970-742.5
.0175X=227.5
X=227.5/.0175
X= 13000 INVWSTED @ 4%
33000-13000=20000 INVESTED @ 2.5%
PROOF
.04*13000+.0225*20000=970
520+450=970
970=970